Automous consumption =100m


Investment spending =100m


Government spending 200milion


Export =150 million


Automous export =100 million


Marginary propensity to consume =1/3


Taxes rate is 1/10


Marginary propensity to import 1/10


YF=2150 million



Questions 1.1 culculate the level of Automous spending in the country


1.2calculate the size of multiplier


1.2 calculate equilibrium level of the income

1
Expert's answer
2021-11-23T10:58:36-0500

1.1

Autonnomous spending = C+I+G+E+X

"=100+100+200+150+100=R650\space million"


1.2

"multiplier=\frac{1}{1-MPC(1-t)+MPM}\\=\frac{1}{(1-\frac{1}{3})(1-\frac{1}{10})+\frac{1}{10}}=0.7"


`1.3

"y=C+I+G+E-M\\Y=100+\frac{1}{3}(Y-\frac{1}{10}Y)+100+200+150-(100+\frac{1}{10}Y)\\Y=550+\frac{3Y}{10}-\frac{Y}{10}\\5Y=550+Y\\Y=R135\space million"